# Towards Multi-Granularity Memory Association and Selection for Long-Term Conversational Agents

### v1

@@ 5 Conclusion
Appendix B Comprehensive analysis of Single-Granularity and Multi-Granularity
B.1 QA Performance
The results in the Table 6 demonstrate the significant advantages of multi-granularity methods over single-granularity approaches. Across all datasets, multi-granularity methods outperform single-granularity methods in key evaluation metrics such as F1, BLEU, ROUGE, and BERTScore. For example, in the LongMemEval-m dataset, the best-performing single-granularity method (Turn-level) achieves an F1 score of 12.26, whereas the multi-granularity method achieves significantly higher F1 scores of 12.51 and 16.85. Similarly, in the LongMTBench+ dataset, the multi-granularity method achieves F1 scores of 37.16 and 41.49, outperforming the best single-granularity method, which scores 36.67. These results highlight the effectiveness of multi-granularity approaches in integrating information from different granularities, providing a more comprehensive understanding and significantly improving QA task performance.
@@ B.1 QA Performance
Table 7: Retrieval Performance.
@@ B.2 Retrieval Performance
The proposed method, MemGAS  further demonstrates its superiority over simple multi-granularity combination methods. In all datasets, MemGAS consistently achieves better results. For example, in the LongMemEval-s dataset, MemGAS achieves Recall@10 of 94.47 and NDCG@10 of 89.96, surpassing the combination method’s scores of 91.91 and 88.03. Similarly, in the LOCOMO-10 dataset, MemGAS achieves Recall@10 of 81.07 and NDCG@10 of 68.24, outperforming the combination method’s scores of 78.70 and 65.64. These improvements demonstrate that MemGAS is not merely aggregating information from multiple granularities but instead employs a more advanced mechanism to capture the interdependencies and complementarities among different granularities. This enables MemGAS to deliver robust and versatile performance, making it highly effective for complex retrieval tasks that require the integration of multi-granularity information.
Appendix C Comparison based on Different Retriever
@@
The results in the Table 8 demonstrate that MemGAS consistently outperforms all other methods across all metrics when evaluated with different base retrievers (MiniLM, MPNet, and Contriever) on the LoCoMo-10 dataset. For the MiniLM base retriever, MemGAS achieves the highest scores in all metrices, outperforming MiniLM, MPC, LLM-Rsum, and SeCom. This indicates that MemGAS is highly effective in leveraging the MiniLM retriever to improve retrieval performance.
@@
Table 8: Retrieval Performance based on different retriever on LoCoMo-10. ’facebook/contriever’, ’sentence-transformers/multi-qa-mpnet-base-cos-v1’, ’sentence-transformers/multi-qa-MiniLM-L6-cos-v1’
@@
MiniLM [43] 42.55 44.19 52.37 49.01 67.98 54.59
MPC [19] 42.30 43.59 51.21 48.08 68.03 54.03
LLM-Rsum [49] 44.76 46.82 54.73 51.64 72.16 57.52
SeCom [33] 45.77 47.25 54.93 51.70 71.15 57.37
@@
MPNet [43] 45.92 47.71 53.98 51.79 68.58 56.88
MPC [19] 45.47 47.35 54.08 51.68 68.28 56.59
LLM-Rsum [49] 49.50 51.15 59.47 56.16 76.64 61.99
SeCom [33] 47.53 49.03 57.05 53.57 70.90 58.57
@@
Contriever [13] 49.90 52.15 58.26 56.29 71.80 60.92
MPC [19] 49.50 51.47 57.45 55.53 71.85 60.47
LLM-Rsum [49] 47.23 48.99 59.01 54.58 74.97 60.07
SeCom [33] 55.24 57.90 64.80 62.36 78.30 66.97
@@
Figure 5: Hyperparameters Analysis on Locomo-10 dataset.
Appendix D Hyperparameter Sensitivity Analysis
@@
Appendix E Comparison on different query types
The radar charts in Figure 6 demonstrate the performance of different methods across diverse query dimensions. Our approach consistently outperforms baseline methods, particularly excelling in multi-hop retrieval on LoCoMo-10 (Figures 6a, 6b) and multi-session on LongMemEval-s (Figures 6c, 6d). This highlights the effectiveness of our framework in addressing both complex reasoning tasks and simpler retrieval-based queries. In multi-hop retrieval task, our method achieves notably higher F1 and RougeL scores, showcasing its ability to retrieve and integrate information across multiple steps. Similarly, in multi-session task, the model effectively captures relevant historical sessions, enabling context-aware and accurate responses. These strengths are driven by the memory association mechanisms that enhance reasoning and knowledge integration.
Figure 6: Comparing the F1 and RougeL socre of different query types on LoCoMo-10 and LongMemEval-s datasets.
MemGAS HippoRAG 2 RAPTOR SeCom LLM-Rsum
Appendix F Cost and Efficiency Analysis
The results in Table 9 clearly demonstrate the exceptional efficiency of MemGAS compared to other baselines on the LongMemEval-s dataset. MemGAS processes only 52.9 M input tokens, which matches the total corpus size, while all other methods require significantly more tokens. For example, HippoRAG 2 processes over 111.1 M input tokens, RAPTOR uses over 62.6 M, and even the relatively efficient LLM-Rsum exceeds MemGAS . Similarly, in terms of output token usage, MemGAS generates only 5.2 M tokens, far fewer than the massive outputs of HippoRAG 2 and SeCom, which produce over 100 M and 70 M tokens, respectively. MemGAS achieves this remarkable efficiency without compromising performance, making it a highly effective and scalable solution for handling large-scale datasets. This demonstrates MemGAS ’s ability to minimize computational costs while maintaining state-of-the-art results.
@@ Appendix G Error Analysis
The error analysis presented in Figure 7 highlights the performance across three datasets: LoCoMo-10, LongMemEval-m, and LongMemEval-s. A key observation is that a significant portion of queries in the LongMemEval datasets lack correct retrieval results, as shown by the high percentage of "Wrong Retrieval + Wrong Generation" (40.6% in LongMemEval-m and 18.6% in LongMemEval-s). Our method effectively identifies cases where the retrieval corpus does not contain information related to the query, responding with "you don’t mention the related information." This approach reduces hallucinations caused by large language models (LLMs) being overly confident and generating fabricated content. For example, in LongMemEval-s, "Correct Retrieval + Correct Generation" accounts for 52.6%, demonstrating the method’s reliability in handling scenarios with relevant information.
@@
Appendix H LLM Prompts
@@
Figures 811 present the key prompt templates used in our multi-stage processing pipeline. Figure 8 defines prompts for multi-granularity information generation, where the model is asked to produce both high-level summaries and fine-grained keyword lists from user-assistant dialogue histories. Figure 9 introduces a filtering prompt that selects only the content relevant to the input question, operating over the structured outputs (session, summary, and keywords) while preserving original token fidelity. Figure 10 provides the instruction for generating final responses, guiding the model to produce concise and coherent answers grounded in the filtered history. Finally, Figure 11 shows the evaluation prompt used to assess answer correctness, following the setup of prior work [50, 58]. This prompt asks GPT-4o to compare model outputs against reference answers and return a binary decision without paraphrasing. Together, these prompts enable modular abstraction, filtering, reasoning, and automatic evaluation across the dialogue memory pipeline.
@@
Figure 10: Prompt for QA, which follows Lu et al. [28], Pan et al. [33]
@@
Figure 11: Prompt for GPT4o-as-Judge (Single), which follows Wu et al. [50], Zheng et al. [58]
Appendix I Case Study
@@
I.1 Case Study on QA Comparison
@@
I.2 Case Study on Multi-granularity Information Generation
@@
[assistant] I’ll definitely look into those options. By the way, do you know how to appraise the value of antique items? ... (Omitted for brevity) [user]:
@@
I.3 Case Study on Multi-granularity Filter
@@
[assistant] Yes, sure. Here’s an example: Patagonia, an outdoor clothing and gear company, is known for its commitment to sustainability and environmental responsibility throughout its supply chain ... (Omitted for brevity) Summary-level information … (Omitted for brevity) Patagonia was highlighted as a prime example, showcasing its commitment to sustainability …(Omitted for brevity) Keyword-level information Patagonia environmentally responsible supply chain
@@
[assistant] That a fantastic step towards reducing your environmental impact! Switching to an eco-friendly ... (Omitted for brevity) Summary-level information The user is considering trying a new grocery delivery service that partners with local farmers…(Omitted for brevity) Keyword-level information meal planning grocery list start small
@@
[assistant] I can give you an example of two companies that prioritize the safety and well-being of their employees like Triumvirate does: 1. Patagonia ... (Omitted for brevity) 2. Southwest Airlines ... (Omitted for brevity) Summary-level information … (Omitted for brevity) to which the AI cited Patagonia and Southwest Airlines as examples of organizations that prioritize employee safety and well-being …(Omitted for brevity) Keyword-level information Triumvirate employee safety tangible improvements
@@
[assistant] That sounds like a great idea! I’m happy to help you with that ... (Omitted for brevity) Summary-level information The user is focused on making eco-friendly changes in their kitchen, specifically looking for sustainable alternatives to kitchen utensils, …(Omitted for brevity) Keyword-level information eco-friendly kitchen makeover sustainable kitchen utensils bamboo utensils
@@
Appendix J Limitations
While our proposed framework demonstrates significant advancements in addressing multi-granularity memory association and adaptive granularity selection, it is not without limitations. There are some key areas where MemGAS can be further improved. Our method depends on LLM-generated summaries and keywords. The quality of memory construction in MemGAS heavily relies on the ability of LLMs to generate accurate and relevant summaries and keywords. Errors or biases in these generated outputs could propagate through the memory system, affecting the accuracy of memory association and retrieval. The scalability of large-scale memory systems needs to be improved. Although MemGAS leverages techniques such as Gaussian Mixture Models and Personalized PageRank to manage and retrieve memory, its computational efficiency may degrade as the memory size grows substantially. This could limit its applicability in scenarios requiring the processing of extensive conversational histories or highly dynamic memory updates.

### v2

@@ Appendix B Comparison of Methods Structure with Baselines
Appendix C Comprehensive Comparison of Single-Granularity and Multi-Granularity
C.1 QA Performance.
The results in the Table 5 demonstrate the significant advantages of multi-granularity methods over single-granularity approaches. Across all datasets, multi-granularity methods outperform single-granularity methods in key evaluation metrics such as F1, BLEU4, ROUGE, and BERTScore. For example, in the LongMemEval-m dataset, the best-performing single-granularity method (Turn-level) achieves an F1 score of 12.26, whereas the multi-granularity method achieves significantly higher F1 scores of 12.51 and 16.85. Similarly, in the LongMTBench+ dataset, the multi-granularity method achieves F1 scores of 37.16 and 41.49, outperforming the best single-granularity method, which scores 36.67. These results highlight the effectiveness of multi-granularity approaches in integrating information from different granularities, providing a more comprehensive understanding and significantly improving QA task performance.
@@ Appendix C Comprehensive Comparison of Single-Granularity and Multi-Granularity
Table 6: Retrieval Performance of Single-Granularity and Multi-Granularity.
@@ C.2 Retrieval Performance
The proposed method, MemGAS  further demonstrates its superiority over simple multi-granularity combination methods. In all datasets, MemGAS consistently achieves better results. For example, in the LongMemEval-s dataset, MemGAS achieves Recall@10 of 94.47 and NDCG@10 of 89.96, surpassing the combination method’s scores of 91.91 and 88.03. Similarly, in the LoCoMo dataset, MemGAS achieves Recall@10 of 81.07 and NDCG@10 of 68.24, outperforming the combination method’s scores of 78.70 and 65.64. These improvements demonstrate that MemGAS is not merely aggregating information from multiple granularities but instead employs a more advanced mechanism to capture the interdependencies and complementarities among different granularities. This enables MemGAS to deliver robust and versatile performance, making it highly effective for complex retrieval tasks that require the integration of multi-granularity information.
C.3 How Multi-granularity Router works?
In this section, we analyze how the proposed Multi-granularity Router works. Our findings reveal that different levels of granularity exhibit distinct preferences on the query type, and our router effectively adapts by selecting the suitable granularity strategy. Table 7 presents the retrieval performance (Recall@3) across various query types and granularities on the LongMemEval-m dataset. The results show that different query types favor different granularities: for instance, temporal-reasoning queries benefit most from session-level granularity, knowledge-update queries achieve better performance with turn-level granularity, and single-session-preference queries perform best with summary-level granularity. Notably, our Granularity Router (MemGAS) adaptively integrates multiple granularities, achieving retrieval performance that closely approaches the upper bound defined by Optimal Selection. This highlights the router’s ability to dynamically identify and apply the most effective granularity for each query type, bridging the gap between fixed granularity approaches and optimal strategies.
Table 7: Comparison of retrieval performance (Recall@3) across query types and granularities, showcasing the effectiveness of the Multi-granularity Router.
Query Type Session Turn Summary Keyword Granularity Router (Ours) Optimal Selection
single-session-assistant 96.43 98.21 96.43 92.86 100.0 100.0
single-session-user 51.56 65.62 54.69 51.56 60.94 78.12
multi-session 32.23 28.1 19.01 13.22 33.88 43.80
knowledge-update 37.5 41.67 37.5 34.72 51.39 72.22
temporal-reasoning 33.86 30.71 33.07 22.83 41.73 51.96
single-session-preference 40.0 33.33 43.0 33.33 46.67 63.33
Appendix D Additional Cost and Efficiency Analysis
D.1 Token Consumption for Memory Construction
In this section, we present the input and output token consumption during memory construction using LLMs (e.g., constructing knowledge graph of HippoRAG2, Hierarchical Tree of RAPTOR, Semantic segment of SeCom and Recursive Summary of RecurSum). The results in Table D.1 clearly demonstrate the exceptional efficiency of MemGAS compared to other baselines on the LongMemEval-s dataset. MemGAS processes only 52.9 M input tokens, which matches the total corpus size, while all other methods require significantly more tokens. For example, HippoRAG 2 processes over 111.1 M input tokens, RAPTOR uses over 62.6 M, and even the relatively efficient RecurSum exceeds MemGAS . Similarly, in terms of output token usage, MemGAS generates only 5.2 M tokens, far fewer than the massive outputs of HippoRAG 2 and SeCom, which produce over 100 M and 70 M tokens, respectively. MemGAS achieves this remarkable efficiency without compromising performance, making it a highly effective and scalable solution for handling large-scale datasets. This demonstrates MemGAS ’s ability to minimize computational costs while maintaining state-of-the-art results.
Extra Memory Cost: We also assessed the additional memory (RAM) cost introduced by our method, as presented in Table 9. On the LongMemeval-s dataset, our multi-granularity approach incorporates summary and keyword memory, resulting in approximately 27MB of extra memory usage—just 10% of the 266MB required for raw memory. This demonstrates that our method imposes minimal storage overhead, further validating its efficiency.
MemGAS HippoRAG 2 RAPTOR SeCom RecurSum
Table 9: Extra Memory (Total Tokens and RAM) Cost on LongMemEval-s datasets.
Original Memory Summary Memory (Extra) Keywords Memory (Extra)
Total Tokens 51.6 milinon 3.14 milinon 2.09 milinon
RAM 266MB 16.2MB 10.8MB
D.2 Comparison at the Same Token Cost
Table 10: Comparison of Methods at Same Token Cost.
Input Tokens Method Latency (s) GPT4o-J
8,000 Contriever 1.81 55.2
Secom 2.13 57.8
HippoRAG2 4.39 57.4
MemGAS (Ours) 2.42 59.8
8,000 Contriever 2.48 56.6
Secom 2.86 58.0
HippoRAG2 4.99 58.2
MemGAS (Ours) 3.15 60.3
103,137 Full history 9.39 50.6
We conduct experiments by fixing the number of input tokens across different methods on the LongMemEval-s dataset. Specifically, we performed over-retrieval and then applied truncation at thresholds of 8,000 and 16,000 input tokens to ensure that all baseline methods have the same input token length. Table 10 are the results of these experiments. The results demonstrate that under the same input token constraints, our proposed method, MemGAS, consistently outperforms baseline approaches in terms of performance (GPT4o-J score) while maintaining competitive latency. Specifically, at both token thresholds, MemGAS achieves the highest GPT4o-J scores (59.8 and 60.3, respectively), surpassing other methods. Although MemGAS incurs slightly higher latency compared to Contriever and Secom, it achieves a significantly better trade-off between efficiency and accuracy. Additionally, the results highlight that using a fixed token cost is more effective than relying on the full history approach, which yields a lower GPT4o-J score of 50.6 despite its much higher latency. This confirms the efficiency and effectiveness of MemGAS in handling long-term memory tasks.
Appendix E Generalization Analysis
E.1 Comparison on Different Retriever
@@
The results in the Table 11 demonstrate that MemGAS consistently outperforms all other methods across all metrics when evaluated with different base retrievers (MiniLM, MPNet, and Contriever) on the LoCoMo dataset. For the MiniLM base retriever, MemGAS achieves the highest scores in all metrices, outperforming MiniLM, MPC, RecurSum, and SeCom. This indicates that MemGAS is highly effective in leveraging the MiniLM retriever to improve retrieval performance.
@@
Table 11: Retrieval Performance based on different retriever on LoCoMo. ’facebook/contriever’, ’sentence-transformers/multi-qa-mpnet-base-cos-v1’, ’sentence-transformers/multi-qa-MiniLM-L6-cos-v1’
@@
MiniLM (Song et al., 2020) 42.55 44.19 52.37 49.01 67.98 54.59
MPC (Lee et al., 2023) 42.30 43.59 51.21 48.08 68.03 54.03
RecurSum (Wang et al., 2025) 44.76 46.82 54.73 51.64 72.16 57.52
SeCom (Pan et al., 2025) 45.77 47.25 54.93 51.70 71.15 57.37
@@
MPNet (Song et al., 2020) 45.92 47.71 53.98 51.79 68.58 56.88
MPC (Lee et al., 2023) 45.47 47.35 54.08 51.68 68.28 56.59
RecurSum (Wang et al., 2025) 49.50 51.15 59.47 56.16 76.64 61.99
SeCom (Pan et al., 2025) 47.53 49.03 57.05 53.57 70.90 58.57
@@
Contriever (Izacard et al., 2021) 49.90 52.15 58.26 56.29 71.80 60.92
MPC (Lee et al., 2023) 49.50 51.47 57.45 55.53 71.85 60.47
RecurSum (Wang et al., 2025) 47.23 48.99 59.01 54.58 74.97 60.07
SeCom (Pan et al., 2025) 55.24 57.90 64.80 62.36 78.30 66.97
@@
E.2 Comparison on Different Generator
We conducted additional experiments using Qwen-3 (8B and 1.7B) as generators, combined with Contriever as the retriever on the LongMemEval-s dataset. Table 12 shows that our proposed MemGASconsistently outperforms the baselines (Contriever and SeCom) across all base generators, including GPT4o-Mini, qwen3-8b, and qwen3-1.7b. Notably, MemGASachieves the highest F1, BLEU4, and Rouge scores, demonstrating its robust ability to both retrieve relevant information and generate high-quality answers. The results also highlight that larger base generators, such as GPT4o-Mini and qwen3-8b, lead to better overall performance, but MemGASmaintains its superiority regardless of the base generator used.
Table 12: QA performance comparison of different models across various base generators (GPT4o-Mini, Qwen-3 8B, and Qwen-3 1.7B) on the LongMemEval-s dataset.
Model GPT4o-J F1 BLEU4 Rouge1 Rouge2 RougeL BertScore
Base Generator: GPT4o-Mini
Contriever (Izacard et al., 2021) 55.40 13.78 2.21 14.46 6.93 12.89 83.70
SeCom (Pan et al., 2025) 56.00 12.95 2.25 13.80 6.09 11.93 83.51
MemGAS 60.20 20.38 4.22 21.05 10.47 19.47 85.21
Base Generator: qwen3-8b
Contriever (Izacard et al., 2021) 50.6 14.76 1.85 15.37 7.44 13.81 83.73
SeCom (Pan et al., 2025) 50.6 14.83 1.89 17.66 8.45 14.08 83.15
MemGAS 51.4 20.57 3.5 21.17 10.15 19.09 85.14
Base Generator: qwen3-1.7b
Contriever (Izacard et al., 2021) 36.4 9.39 1.3 9.92 4.37 8.8 82.62
SeCom (Pan et al., 2025) 36.0 8.5 1.07 9.26 3.66 7.91 82.26
MemGAS 38.0 16.79 3.6 17.58 7.78 16.19 84.84
E.3 Comparison on Different Query Types
The radar charts in Figure 5 demonstrate the performance of different methods across diverse query dimensions. Our approach consistently outperforms baseline methods, particularly excelling in multi-hop retrieval on LoCoMo (Figures 5a, 5b) and multi-session on LongMemEval-s (Figures 5c, 5d). This highlights the effectiveness of our framework in addressing both complex reasoning tasks and simpler retrieval-based queries. In multi-hop retrieval task, our method achieves notably higher F1 and RougeL scores, showcasing its ability to retrieve and integrate information across multiple steps. Similarly, in multi-session task, the model effectively captures relevant historical sessions, enabling context-aware and accurate responses. These strengths are driven by the memory association mechanisms that enhance reasoning and knowledge integration.
Figure 5: Comparing the F1 and RougeL socre of different query types on LoCoMo and LongMemEval-s datasets.
Figure 6: Hyperparameters Analysis on LoCoMo dataset.
Appendix F Hyperparameter Sensitivity Analysis
@@ Appendix G Error Analysis
The error analysis presented in Figure 7 highlights the performance across three datasets: LoCoMo, LongMemEval-m, and LongMemEval-s. A key observation is that a significant portion of queries in the LongMemEval datasets lack correct retrieval results, as shown by the high percentage of "Wrong Retrieval + Wrong Generation" (40.6% in LongMemEval-m and 18.6% in LongMemEval-s). Our method effectively identifies cases where the retrieval corpus does not contain information related to the query, responding with "you don’t mention the related information." This approach reduces hallucinations caused by large language models (LLMs) being overly confident and generating fabricated content. For example, in LongMemEval-s, "Correct Retrieval + Correct Generation" accounts for 52.6%, demonstrating the method’s reliability in handling scenarios with relevant information.
@@
Appendix H Theoretical Analysis
We provide concise guarantees for two building blocks of MemGAS: (i) GMM-based accept/reject association, and (ii) multi-granularity scoring with entropy-driven routing. All results are stated under standard, verifiable assumptions with compact proof sketches.
Notation.
For a query $q$ and memory index $i\in\{1,\dots,n\}$, MemGAS maintains four granularities $g\in\{S\ (\text{session}),\,T\ (\text{turn}),\,U\ (\text{summary}),\,K\ (\text{keyword})\}$ with similarity scores $s_{i}^{g}=\mathrm{sim}(q,M_{i}^{g})$. For each $g$, define the softmax distribution and entropy
$p_{g}(i)\;=\;\frac{\exp(s_{i}^{g}/\lambda)}{\sum_{j=1}^{n}\exp(s_{j}^{g}/\lambda)},\qquad H_{g}\;=\;-\sum_{i=1}^{n}p_{g}(i)\log p_{g}(i).$
The router weight is $w_{g}=H_{g}^{-1}/\sum_{g^{\prime}}H_{g^{\prime}}^{-1}$, and the initial aggregate score is $\mathrm{score}_{i}=\sum_{g}w_{g}\,s_{i}^{g}$ with normalized vector $s=(\mathrm{score}_{i})_{i}/\sum_{j}\mathrm{score}_{j}$. Let $G=(V,E)$ denote the association graph induced by the accept/reject rule. Let $R(q)\subseteq\{1,\dots,n\}$ be the relevant set, Recall@$K$ the top-$K$ recall, and nDCG the ranking metric. For each granularity $g$, let $TopK_{g}(q)$ denote the indices of its top-$K$ scored items.
Assumptions.
(A1) (Sub-Gaussian separability) For each $g$, $s_{i}^{g}|y_{i}=1$ and $s_{i}^{g}|y_{i}=0$ are sub-Gaussian with means $\mu_{g}^{+}>\mu_{g}^{-}$ and common scale $\sigma_{g}$; denote $\Delta_{g}=\mu_{g}^{+}-\mu_{g}^{-}>0$.
H.1 GMM Accept/Reject Association (Clean Links)
Intuition. We want the association graph to connect new memories mostly to truly related ones. If the relevance/non-relevance score distributions are well-separated, a simple threshold already yields exponentially small mistakes; a fitted GMM recovers (approximately) this threshold in practice.
Proposition 1 (Exponentially small mis-link rate)
Under A1 (drop index $g$ for brevity), the mid-threshold $\tau^{\star}=(\mu^{+}+\mu^{-})/2$ satisfies
[…]
Moreover, for identifiable mixtures, GMM consistently learns $(\mu^{\pm},\sigma)$, attaining the same exponential error order. (Sketch) Apply sub-Gaussian tail bounds at $\mu^{\pm}\mp\Delta/2$ and union bound.
Corollary 1 (Few false edges in expectation)
Let $\eta_{g}\leq 2\exp(-\Delta_{g}^{2}/(8\sigma_{g}^{2}))$ be the accept/reject error rate at granularity $g$, and $d_{g}$ the number of candidate neighbors. Then the expected number of false edges per insertion satisfies $\mathbb{E}[|E_{\mathrm{false}}|]=\mathcal{O}(\sum_{g}d_{g}\eta_{g})$ (exponentially small in $\Delta_{g}$).
Consequence. With larger separation $\Delta_{g}$, the graph stays “clean” over time, which is crucial for subsequent retrieval and ranking stages.
H.2 Entropy Router (Choose Confident Granularities)
Intuition. Each granularity produces a relevance distribution $p_{g}$ over memories. If this distribution is sharp (low entropy), that granularity is more decisive for the current query. Weighting granularities by inverse entropy therefore favors the most confident signals without hand-tuning.
Lemma 1 (Lower entropy $\Rightarrow$ higher confidence)
If the distribution $p_{g}$ becomes more concentrated on its top-1 item (larger top-1 margin), its entropy $H_{g}$ decreases. (Sketch) Entropy is Schur-concave: more concentrated distributions have lower entropy.
Proposition 2 (Inverse-entropy weighting is simple and robust)
The rule $w_{g}\propto 1/H_{g}$ is (i) monotone in confidence (Lemma 1), (ii) symmetric across granularities, and (iii) scale-free under common rescaling of $\{H_{g}\}$. (Sketch) These invariances restrict $w_{g}$ to power laws $H_{g}^{-\beta}$; $\beta=1$ is the simplest scale-free choice.
Theorem 1 (Multi-granularity candidate-pool coverage is never worse)
Let the multi-granularity candidate pool be $C_{\mathrm{multi}}=\bigcup_{g}\textsf{TopK}_{g}(q)$. Its coverage of relevant items is
$\frac{|C_{\mathrm{multi}}\cap R(q)|}{|R(q)|}\;\geq\;\max_{g}\mathrm{Recall@}K_{(g)},$
$\text{and under weak dependence}\;\;\frac{|C_{\mathrm{multi}}\cap R(q)|}{|R(q)|}\;\gtrsim\;1-\prod_{g}\big(1-\mathrm{Recall@}K_{(g)}\big).$
(Sketch) A union cannot cover fewer relevant items than any constituent set; inclusion–exclusion yields the product bound under weak dependence. MemGAS approximates this union via weighted aggregation.
Consequence. Adaptive combination cannot underperform the best single granularity for recall, and typically exceeds it when granularities are complementary (e.g., summaries remove noise; turns capture details).
Discussion. Why does MemGAS beat single/fixed granularity? Under A1 and for suitable $K$: (i) the multi-granularity candidate pool, guided by $w_{g}\propto 1/H_{g}$, achieves coverage at least as high as the best single granularity (Thm. 1) and typically higher when the views are complementary; (ii) inverse-entropy routing (Prop. 2) adaptively emphasizes the most informative granularity per query, improving the quality of the aggregated scores without hand-tuning; and (iii) the GMM accept/reject mechanism yields exponentially few false associations (Prop. 1 and its corollary), reducing noise and spurious ties. Together, these effects raise Recall@K of the candidate pool and typically improve nDCG once any standard downstream ranker is applied. Any additional retrieval module (e.g., graph propagation) is orthogonal and can be plugged in as desired, but it is not a contribution of our method.
Appendix I LLM Prompts Design
@@
Figures 811 present the key prompt templates used in our multi-stage processing pipeline. Figure 8 defines prompts for multi-granularity information generation, where the model is asked to produce both high-level summaries and fine-grained keyword lists from user-assistant dialogue histories. Figure 9 introduces a filtering prompt that selects only the content relevant to the input question, operating over the structured outputs (session, summary, and keywords) while preserving original token fidelity. Figure 10 provides the instruction for generating final responses, guiding the model to produce concise and coherent answers grounded in the filtered history. Finally, Figure 11 shows the evaluation prompt used to assess answer correctness, following the setup of prior work (Wu et al., 2025a; Zheng et al., 2023). This prompt asks GPT-4o to compare model outputs against reference answers and return a binary decision without paraphrasing. Together, these prompts enable modular abstraction, filtering, reasoning, and automatic evaluation across the dialogue memory pipeline.
@@
Figure 10: Prompt for QA, which follows Lu et al. (2023); Pan et al. (2025)
@@
Figure 11: Prompt for GPT4o-as-Judge (Single), which follows Wu et al. (2025a); Zheng et al. (2023)
Appendix J Case Study
@@
J.1 Case Study on QA Comparison
@@
J.2 Case Study on Multi-granularity Information Generation
@@
[assistant] I’$l$ definitely look into those options. By the way, do you know how to appraise the value of antique items? ... (Omitted for brevity) [user]:
@@
J.3 Case Study on Multi-granularity Filter
@@
[assistant] Yes, sure. Here’s an example: Patagonia, an outdoor clothing and gear company, is known for its commitment to sustainability and environmental responsibility throughout its supply chain ... (Omitted for brevity) Summary-level information (Omitted for brevity) Patagonia was highlighted as a prime example, showcasing its commitment to sustainability …(Omitted for brevity) Keyword-level information Patagonia environmentally responsible supply chain
@@
[assistant] That’s a fantastic step towards reducing your environmental impact! Switching to an eco-friendly ... (Omitted for brevity) Summary-level information The user is considering trying a new grocery delivery service that partners with local farmers …(Omitted for brevity) Keyword-level information meal planning grocery list start small
@@
[assistant] I can give you an example of two companies that prioritize the safety and well-being of their employees like Triumvirate does: 1. Patagonia ... (Omitted for brevity) 2. Southwest Airlines ... (Omitted for brevity) Summary-level information (Omitted for brevity) to which the AI cited Patagonia and Southwest Airlines as examples of organizations that prioritize employee safety and well-being …(Omitted for brevity) Keyword-level information Triumvirate employee safety
@@
[assistant] That sounds like a great idea! I’m happy to help you with that ... (Omitted for brevity) Summary-level information The user is focused on making eco-friendly changes in their kitchen, specifically looking for sustainable alternatives to kitchen utensils, …(Omitted for brevity) Keyword-level information eco-friendly kitchen makeover sustainable kitchen utensils